OpenStudy (anonymous):

Density of Cu is 8.92 g/cc.Density of water is 10 raised to the power of 3 cube i.e.,10 cube3kg/m cube3.Determine the relative density of Cu. Ramya

OpenStudy (anonymous):

Hi ash2326. can you also help me with the question stated below ? swecho: Density of Cu is 8.92 g/cc.Density of water is 10 raised to the power of 3 cube i.e.,10 cube3kg/m cube3.Determine the relative density of Cu… 1 viewing

OpenStudy (fretje):

Density is represented in physics often by greek letter Rho and defined as the mass over volume (Rho = m/V) . cc is the abreviation for cubic centimeter, which is a unit of volume.Thus we write. Rho Cu = 8,92 gram/cm³ = 8,92 g/cm³ * 1kg/1000g * m³/1000000cm³ = 8920 kg/m³ Rho Water = 1000kg/m³ Now we can compare both matters, when we name the relative density of copper d Cu: d Cu = Rho Cu/Rho Water = (8920kg/m³) / (1000kg/m³) = 8,92

OpenStudy (anonymous):

thank u so much,fretje!!